nts 4 y (m) time 74s 25 15 10 OS 10 15 20 25 30 35 The an mation shows a bail which has been kicked upward at an angle. Ran the animation to watch the motion of the bail Tthe ball does not show when you hit play, try cikking hete to start the animation et by the ballonce par second the animation can aso be passed and moved forward in single frame mode ifyoud lke The cursor can be used to resd the cordinates of a postion in the grid by holding dew presume that x s positivt to the r ght a d、is positive upwards Note that this bala net be g keke en Earth. Do not exped an accete aten ers o mns use this a to ano e me foi me g que- mato components at is the maximum height which the bailreaches? 20.12 m are correct What s the horizontal component of the initial velocity of the bal? 7/10 Previous Tries Submt Answer Incorrect Tries Tno What is the vertical component of the initial velocity of the bail? 10 .88 m/e You are correct recept no a 152 9209 0) Previus Tries What s the vertical component of the acceleration of the bail?-2.94 m/s 2
Answer #1

Time of flight = 2Vy/ay

here Vy is vertical component of initial velocity

ay is vertical component of acceleration.

so T = 2*10.88/2.94 = 7.4 sec

horizontal range = VxT

here Vx is horizontal component of velocity.

from graph horizontal range = 34 m

34 = Vx*7.4

Vx = 4.59 m/sec